OpenStudy (anonymous):

Two variables p and q are related by the equation p=3q+2/q . Given q increases at a constant rate of 4 units per second, find the positive value of q at the instant p is increasing at a rate of 10 units per second . Just to double check my working @ganeshie8 @jhonyy9

jhonyy9 (jhonyy9):

do you think p = (3p +2)/q or p = 3p + (2/q) how is correct ?

ganeshie8 (ganeshie8):

Hey!

OpenStudy (anonymous):

\[p =3q+\frac{ 2 }{ q }\]

OpenStudy (anonymous):

Hey @ganeshie8 :)

jhonyy9 (jhonyy9):

so than p = (3q^2 +2)/q - q increase 4 unit/sec. so this mean that p increase p = (3*16 +2)/4 = 25/2 = 12,5 unit/sec - but you need to calculi the positive value of q at the instant p is increasing at a rate of 10 unit per sec. so than you can writing 10 = (3q^2 +2)/q - solve this quadratic for positive value of q and will get the right answer to your exercise sure - hope so much that this will help you and that this is understandably - good luck

OpenStudy (anonymous):

i did like this...Hope i'm right..\[p=3q+\frac{ q }{ 2 }\]\[p=3q+2q^{-1}\]\[\frac{ dp }{ dq }=3-\frac{ 2 }{ q^2 }\]\[\frac{ dp }{ dt }=\frac{ dp }{ dq }\times \frac{ dq }{ dt }\]\[10=(3-\frac{ 2 }{ q })\times4\]\[\frac{ 5 }{ 2 }=3-\frac{ 2 }{ q^2 }\]\[\frac{ 2 }{ q^2 }=\frac{ 1 }{ 2 }\]\[q^2=4\]\[q=\pm \sqrt{4}\]\[q=2,\cancel{q=-2}\]

jhonyy9 (jhonyy9):

so for you can being secure in your result you can making the prove for this q=2 so substituting in place of q inside this formula p = 3q +2/q 10 = 3*2 +2/2 10 = 6+1 10 = 7 - so this mean that not is right ATTENTION !!!

jhonyy9 (jhonyy9):

@MARC above you have wrote p= 3q + (q/2) but in the text of your exercise there you have wrote p = 3q + (2/q) - can you deciding please how is exactly sure correct this text of your exercise ,please ???

OpenStudy (anonymous):

i'm sry @jhonyy9 it was an typo error... but the equation is\[p=3q+\frac{ 2 }{ q }\]

OpenStudy (anonymous):

@jhonyy9

jhonyy9 (jhonyy9):

@MARC so than p = 3q +2/q is right that you can correct your calculi because yu have doit them with p = 3q + q/2 please check it and correct your calculi for you can giving the right answer

OpenStudy (anonymous):

alright,let me correct it back again :)

OpenStudy (anonymous):

This is my working..\[p=3q+\frac{ 2 }{ q }\]\[p=3q+2q^{-1}\]\[\frac{ dp }{ dq }=3-\frac{ 2 }{ q^2 } \]\[\frac{ dp }{ dt }=\frac{ dp }{ dq }\times \frac{ dq }{ dt } \]\[10=(3-\frac{ 2 }{ q })\times4 \]\[\frac{ 5 }{ 2 }=3-\frac{ 2 }{ q^2 } \]\[\frac{ 2 }{ q^2 }=\frac{ 1 }{ 2 } \]\[q^2=4 \]\[q=\pm \sqrt{4} \]\[q=2,\cancel{q=-2} \]

jhonyy9 (jhonyy9):

@MARC_ so check it please above i have doit the prove of q=2 and you dont give p=10 so what mean that this is something wrong in your calculi

OpenStudy (anonymous):

is it q=3.120 and q=0.214 ? @jhonyy9

jhonyy9 (jhonyy9):

so than you need make the prove of these results and check that you will got p=10 ok ? check for what result of q will get p=10 sure ?

jhonyy9 (jhonyy9):

p = 3*3 + 2/3 10 = 29/3 what not is right sure --- yes ?

OpenStudy (anonymous):

\[10\neq \frac{ 29 }{ 3 }\]

jhonyy9 (jhonyy9):

yes so this mean that q=3 not is right root of equation

OpenStudy (anonymous):

q=0.214 is also not the right root of equation

jhonyy9 (jhonyy9):

than this mean that your calculi are wrong

OpenStudy (anonymous):

can u show me your working pls.. :) so that i can know where i'm wrong

jhonyy9 (jhonyy9):

10 = 3q + 2/q so this a very easy quadratic you dont know solve it right ??? --- why ?

jhonyy9 (jhonyy9):

how you have got it here ?

OpenStudy (anonymous):

\[3q^2+2=10q\]

jhonyy9 (jhonyy9):

yes

OpenStudy (anonymous):

\[3q^2-10q+2=0\]

jhonyy9 (jhonyy9):

yes this the quadratic what you need to solve it right

OpenStudy (anonymous):

\[q=3.120~and~q=0.214\]

jhonyy9 (jhonyy9):

sure ?

OpenStudy (anonymous):

That's the value i get from the calculator but it seems that i'm not getting p=10 when i substitute the value of q into the equation p=3q+(2/q)..

jhonyy9 (jhonyy9):

why ? this mean that there are some mistaque ?

jhonyy9 (jhonyy9):

3q^2 -10q +2 =0 discriminant = ?

OpenStudy (anonymous):

is it \[b^2-4ac>0\]

jhonyy9 (jhonyy9):

please you write here your calculi for i can checking where are you doit wrong

OpenStudy (anonymous):

sry,i don't get u r u asking me solve the equation using the b^2-4ac>0 ?

jhonyy9 (jhonyy9):

yes - you get for discriminant 76 - yes so what not is a perfect square

OpenStudy (anonymous):

i'm not sure can u tell me @jhonyy9

jhonyy9 (jhonyy9):

3q^2 -10q +2 = 0 D = 76 q1,2 = ? hope you know formula to sole it - yes ?

jhonyy9 (jhonyy9):

to solve it

OpenStudy (anonymous):

what formula ?

jhonyy9 (jhonyy9):

formula how you calculi the roots of a quadratic with discriminant greater than zero

OpenStudy (anonymous):

\[b^2-4ac>0\]\[(-10)^2-4(3)(2)>0\]\[76>0\]

jhonyy9 (jhonyy9):

yes but this is how you calculi the discriminant so what mean that the discriminant is greater than zero ?

jhonyy9 (jhonyy9):

so this mean that this quadratic has two different real roots - yes ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so continue calculi these roots of

jhonyy9 (jhonyy9):

formula for ?

OpenStudy (anonymous):

\[3q^2-10q+2>0\]

jhonyy9 (jhonyy9):

q1,2 = (-b+/- sqrt(discriminant) / 2a yes ?

jhonyy9 (jhonyy9):

you dont know this formula ?

OpenStudy (anonymous):

u mean the quadratic formula

jhonyy9 (jhonyy9):

what i have wrote there above now

jhonyy9 (jhonyy9):

for q1,2

OpenStudy (anonymous):

\[q=\frac{ -b~\pm \sqrt{b^2-4ac} }{ 2a }\]\[q=\frac{ -(-10)\pm \sqrt{76} }{ 2(3) }\]\[q=3.120~and~q=0.214\]

jhonyy9 (jhonyy9):

so than you substitute 3,120 you get p=10 ?

OpenStudy (anonymous):

i got p=10.00102564

jhonyy9 (jhonyy9):

so this will be right ?

jhonyy9 (jhonyy9):

but now i have a different idea for you can solve it

jhonyy9 (jhonyy9):

using the rule of 3 simple do you know it ?

OpenStudy (anonymous):

no :/

jhonyy9 (jhonyy9):

when q increase 4 unit/sec so than you make the calculi you get p= 12,5 unit/sec q p 4 u/sec 12,5 u/sec x 10 u/sec ------------------------------------- x = (10*4)/12,5 = 3,2 u/sec x = 3,2 u/sec - do you like this method ? do you understand it ?

jhonyy9 (jhonyy9):

so but i think that your result of 3120 will be more exactly like this 3,2 yes ?

OpenStudy (anonymous):

but when u put q=3.2 into the equation p=3q+(2/q) it does not get p=10.

jhonyy9 (jhonyy9):

yes i understand you but above how you have got this 10,0001... result for p for what substituted root of q ?

jhonyy9 (jhonyy9):

i think 3,120 - yes ?

OpenStudy (mathmale):

This has been a long conversation! MARC_ : What do you still need to know, at this point?

OpenStudy (anonymous):

when u substitue q=0.214 u will also get p=10 when u round off but the question say that find the positve value of q.. so q can't be two values right?

jhonyy9 (jhonyy9):

@mathmale your opinion please ?

OpenStudy (anonymous):

The topic for this question is differentiation

OpenStudy (zarkon):

the answer is 2. as you posted above

jhonyy9 (jhonyy9):

thank you @ganeshie8 so i have a question please - than you see above where Marc have got q=2 have made these calculi for equation when p= 3q + q/2 - after he have said that correct this equation wann being p = 3q + 2/q - what is your opinion about this ,please ?

OpenStudy (anonymous):

Thnx @Zarkon

hartnn (hartnn):

@jhonyy9 "p = 3q +2/q 10 = 3*2 +2/2" here, p is not given as 10, dp/dt is = 10. thats what confused you. MARC's work is fine.

OpenStudy (anonymous):

Thnx @hartnn